OpenStudy (anonymous):

Please help! Thank you! Evaluate f(2), f(-3,2), and f(a) f(t) = -16t(to the second power) + 7.8t + 12

OpenStudy (anonymous):

what is f(-3,2) ??

OpenStudy (anonymous):

Sorry! I meant -3.2.

OpenStudy (anonymous):

I see ;)

OpenStudy (anonymous):

for the first one we insert t=2: \[f(t) = -16(2)^2 + 7.8\times(2) + 12\]Because exponents have a higher priority than multiplication, the (2) must first be squared before being multiplied with 16

OpenStudy (anonymous):

we have three parts of the equation here, separated by + can you calculate the value for the first part?

OpenStudy (anonymous):

-91.6

OpenStudy (anonymous):

-16(2)^2 = -91.6 ?